# FlorDB: Multiversion Hindsight Logging for Continuous Training

### 06 Feb 2024 v2

@@ 6. Evaluation
This evaluation aims to demonstrate the key strengths of FlorDB: its ability to efficiently handle multiversion hindsight logging and its potential for significant speed-ups during replay. We showcase FlorDB’s versatility by applying it to a diverse range of model architectures, as detailed in Table 1. This table summarizes the models’ characteristics, evaluation metrics, and real-world applications. Building upon our prior work that established Flor’s low-overhead recording and parallel replay capabilities for single versions (Garcia et al., 2020), this study focuses on FlorDB’s ability to generalize and scale effectively across multiple versions. Our evaluation delves into two key areas:
  1. (1)
    Scalability: We assess FlorDB’s efficiency in handling growing numbers of model versions, ensuring linear scaling without performance degradation or cross-version interference.
  2. (2)
    Performance: We assess FlorDB’s techniques for accelerating hindsight logging replay. This includes accurate cost estimation for query refinement, capitalizing on cross-version parallelism for speed-ups from scale-out, and efficient utilization of within-version parallelism via checkpoint-based training resume.
@@ 6.1. Efficient Linear Scaling across Versions
Multiversion hindsight logging requires more computation than prior work on a single version, so we wish to demonstrate the scalability of FlorDB, and the absence of cross-version interference. We find that FlorDB scales efficiently and linearly with the number of versions.
6.1.1. Generalizing to $N$ Versions
We tested FlorDB across various model architectures, as shown in 8. Results show that the time complexity of FlorDB grows linearly with the number of versions (y-axis in log scale). This is a significant result, indicating the absence of interaction effects or accumulating overhead, even when dealing with big models or large datasets.
@@ 6.1.2. Interactive Response Times
A standout observation is the consistency in replay times across the different models. Despite the variability in the prefix phase due to variations between data preparation and featurization pipelines, most hindsight logging queries can be answered in near-real time (see brown lines in 8): even for more time-intensive record phases, there are multiple expedient replay options available to users.
@@ 6.1.3. Storage Requirements
9 shows the Git repository, despite frequent commits for dozens of versions, remains remarkably compact, under 5MB. However, checkpoints for large models stored on the shared drive can easily exceed hundreds of gigabytes (brown bars in the left pane). To address this, we propose “Kahneman Sampling” inspired by Kahneman (2011). This approach stores only two checkpoints per version: the “peak” checkpoint based on a chosen metric and the “end” checkpoint at training completion (olivegreen bars in the left pane). This minimal set surprisingly supports all replay-prefix and replay-suffix queries and enables restoring other checkpoints in half the time with 2x parallelism.
Figure 10. Boxplot distribution of errors between actual and expected execution times across different components of a replay execution. Components exhibit low errors on average, and estimates are highly correlated with projected runtime.
@@ 6.2. Speed-up through Selectivity & Scale-out
Estimating a replay query’s runtime is crucial for dynamic ML workloads where time-to-insight is paramount. This helps developers avoid prohibitively long execution times and efficiently utilize resources. As described in 5.2 and depicted in 10, we leverage record profiling statistics to provide highly accurate replay cost estimations (error $\leq 5\%$). These estimates empower developers to refine their queries until further filtering is no longer desirable. At that point, parallelism comes into play. Our prior work on Flor capitalizes on the embarrassingly parallel nature of training replay from checkpoints (Garcia et al., 2020). By scaling out operations across epochs (and also versions), Flor significantly speeds up replay. Therefore, our approach to reducing time-to-insight is two-pronged: first, accurate cost estimation for informed query refinement, and then, when further refinement is no longer desirable, efficient leverage of Flor’s checkpoint-resume parallelism for speed-up through scale-out.
6.2.1. Near-Ideal Parallelism
11 (top) depicts the scale-out ceiling, which represents the time, in minutes, required to run a single epoch of each experiment when parallelism is maximized: that is, when the degree of parallelism is set to equal the number of epochs. Prior studies confirm that Flor approaches ideal parallelism on a single version (Garcia et al., 2020), and FlorDB further leverages embarrassing parallelism from replay across versions.
6.2.2. Cost and Performance Considerations
Parallelism can accelerate computation, but it is also important to evaluate the associated costs. The estimated cost per epoch for various experiments on both CPU and GPU platforms is shown in 11 (bottom). Estimated cost is based on AWS EC2 instance rental fees, with and without a GPU, for the required execution time. While GPUs offer faster computation, they are associated with higher costs.
@@ 8. Conclusion
This paper introduces FlorDB, a system designed to address the unique challenges faced by machine learning engineers (MLEs) in managing the iterative model development process. FlorDB’s approach to multiversion hindsight logging, a record-replay technique, allows MLEs to add logging statements post-hoc, thereby enabling MLEs to “query the past.” FlorDB’s key contributions include a unified relational model for querying log results, automatic propagation of new logging statements across versions as part of acquisitional query processing, and a highly-accurate empirical cost predictor for replay query refinement. These features streamline the ML experimentation process, enabling more efficient analysis and faster iteration. We evaluate the system’s performance across various computer vision and NLP benchmarks, demonstrating the scalability of FlorDB when handling multiversion hindsight logging.

### 02 Mar 2024 v3

@@ 6. Evaluation
This evaluation aims to demonstrate the key strengths of FlorDB: its ability to efficiently handle multiversion hindsight logging and its potential for significant speed-ups during replay. We showcase FlorDB’s versatility by applying it to a diverse range of model architectures, as detailed in Table 1. This table summarizes each model’s characteristics, evaluation metrics, and real-world applications. Building upon our prior work that established Flor’s low-overhead recording and parallel replay capabilities for single versions (Garcia et al., 2020), this study focuses on FlorDB’s ability to generalize and scale effectively across multiple versions. Our evaluation delves into two key areas: scalability and responsiveness. FlorDB’s ability to handle growing numbers of model versions is assessed, ensuring linear scaling without performance degradation or cross-version interference. Experiments were run on our own server with 4 CPU cores (11th Gen Intel Core i7), 32 GB of RAM, 1 TB SSD, and a GeForce RTX 3070 GPU with 8GB GDDR6.
@@ 6.1. Efficient Linear Scaling across Versions
To assess FlorDB’s ability to manage multiversion hindsight logging without cross-version interference, we evaluated its performance across different model architectures. For each model architecture, we measured the runtime for the following modes of execution (as defined in 5.2):
  • record: First run of model training; generates checkpoints.
  • replay-prefix: query runs a prefix scan over all versions.
  • replay-suffix: query runs a suffix scan over all versions.
  • replay-validation: query runs a validation scan over all versions.
Our findings, depicted in 9, reveal that FlorDB’s time complexity scales linearly with an increasing number of versions, as demonstrated by the log-scale on the y-axis. This evidence of linear scalability, devoid of interaction effects or incremental overhead, validates FlorDB’s ability to handle both large models and big datasets.
@@ 6.1.1. Interactive Response Times
In evaluating the impact of different replay operations on system response times (see orange lines in 9), the replay-prefix operation maintains the fastest processing times, consistently delivering results in under 10 seconds. Even as the number of versions increases, the processing time for replay-prefix remains the most stable, ensuring a responsive user experience. In contrast, replay-suffix and replay-validation operations exhibit longer processing times as the version count rises, with replay-validation times increasing more steeply. One factor is the amount of time it takes to evaluate the model. While replay-suffix may still offer interactive speeds at lower version counts, replay-validation quickly surpasses the interactive threshold: in scenarios where immediate feedback is crucial, users can refine the selectivity of their queries for faster response times (as discussed in 5.2).
@@ 6.1.2. Storage Requirements
Despite frequent commits across numerous versions, Git repositories remain remarkably compact, typically under 5MB (see right pane of 10). In contrast, checkpoints for large models stored on shared storage can easily consume hundreds of gigabytes (see orange bars in the left pane). To address potential local storage depletion, FlorDB’s default approach is to spool least-recently-used checkpoints to a cloud object store like S3444Storing 100 GB of data in S3 costs approximately \$2.30 a month.. If this is not feasible, FlorDB may retain a subset of checkpoint per version—as few as just two (see green bars in 10)—evicting the rest. FlorDB retains checkpoints that are evenly spaced across iterations. This ensures balanced work partitions and avoids stragglers that would hinder replay parallelism. Importantly, even retaining just two checkpoints (one mid-run, one at completion) offers significant benefits: i) the final checkpoint enables fast restores for replay-prefix and replay-suffix queries, and the mid-run checkpoint enables 2x faster replay execution compared to the original run.
@@ 6.2. Speed-up through Selectivity & Scale-out
As discussed in the previous section, multiversion hindsight logging queries may be long-running tasks. In order to maintain interactive response times, it may be desirable to refine the query to be more selective (Section 5.2), or to allocate more resources (e.g., GPUs) and run the query in parallel. In some cases, a developer may be willing to pay the cost of generating extensive hindsight log messages from the inner loop of their training script. In this case each iteration of the inner loop must be replayed from checkpoint, and logs generated; a compute- and data-intensive task. In these cases, parallelism comes into play: because each loop iteration is based on a separate checkpoint, we can replay as many loop iterations as possible at once with embarrassing parallelism. We refer the reader to the experimental results in our earlier work measuring this effect (Garcia et al., 2020). To address the potential for long query times and the desire for selectivity or resource scaling, let’s now examine the bottlenecks involved and the cost/performance trade-offs associated with resource use.
6.2.1. Bottleneck Analysis — length-1 range scan
11 (top) depicts the runtime of a range scan over a single epoch. Assuming unbounded resources (e.g. in an idealized cloud), every version would be replayed in its own machine. In this ideal scenario, a query’s bottleneck would be the time it takes for a single Flor replay operation to finish. Replay-prefix, replay-suffix, and replay-validation would each finish in about 100 seconds or less (9). In contrast, the cost of a range scan, even for a single epoch, is higher. This operation requires stepping into the nested training loop, performing a forward pass over the neural network, and back-propagating gradients. Consequently, the range scan emerges as a bottleneck operator. In 11 we evaluate its runtime on a single epoch; running times for more epochs or versions can be linearly extrapolated. This is due to the ideal parallelism of replay-from-checkpoint (confirmed by prior studies (Garcia et al., 2020)) and the embarrassing parallelism of replay across versions.
6.2.2. Cost and Performance Trade-offs
Our bottleneck analysis assumed an ideal of unbounded resources. While parallelism can significantly accelerate computation, it comes with increased resource consumption and financial costs. 11 (bottom) compares the estimated cost per epoch across CPU and GPU platforms. Costs are based on AWS EC2 instance rental fees, with and without GPUs, factoring in execution time. GPUs provide order-of-magnitude superior computational speed but incur order-of-magnitude higher expenses. CPUs have the added benefit that they can be elastically allocated en masse. This can improve throughput across many versions or epochs, but as the bottleneck analysis indicates, there are limits to response times when range scans are used.
@@ 8. Conclusion
This paper introduces FlorDB, a system designed to address the unique challenges faced by machine learning engineers (MLEs) in managing the iterative, data-rich model development process. FlorDB’s approach to multiversion hindsight logging, a record-replay technique, allows MLEs to add logging statements post-hoc, thereby enabling MLEs to “query the past.” FlorDB’s key contributions include a unified relational model for querying log results, automatic propagation of new logging statements across versions as part of acquisitional query processing, and a highly-accurate empirical cost predictor for replay query refinement. These features streamline the ML experimentation process, enabling more efficient analysis and faster iteration. We evaluate the system’s performance across various computer vision and NLP benchmarks, demonstrating the scalability of FlorDB across multiple versions.
Acknowledgements.
This work is supported in part by National Science Foundation CISE Expeditions Award CCF-1730628, IIS-1955488, IIS-2027575, DOE award DE-SC0016260, ARO award W911NF2110339, and ONR award N00014-21-1-2724. Rolando Garcia is supported in part by a grant from Dell Inc. We also thank Robert Lincourt, Matei Zaharia, and Bobby Yan for helpful comments and feedback.